OpenStudy (olivya999):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> A) 3.0° B) 6.0° C) -7.0° D) 16.0°

OpenStudy (olivya999):

@pooja195 @Mehek14 @TheSmartOne

OpenStudy (zpupster):

\[u \cdot v = \left| \left| u \right| \right|\left| \left| v \right| \right|\cos \Theta\]

OpenStudy (zpupster):

so u = (2*3) and v= (-4(-8)= 6+32=38 ||u||= sqrt(2^2 + 3^2) ||v|| = sqrt((-4)^2 + (-8)^2 so 38/(sqrt(2^2 + 3^2))(sqrt((-4^2) + (-8)^2) = COS(THETA)

OpenStudy (zpupster):

theta = 38/(sqrt(2^2 + 3^2))(sqrt((-4^2) + (-8)^2)